Find f(x-6) for f(x)= -2x^2+9x+1
Answer by stanbon(75887)

Find f(x-6) for f(x)= -2x^2+9x+1
----------------
f(x-6) = -2(x-6)^2 + 9(x-6) + 1
------
= -2(x^2-12x+36) + 9x-54 + 1
-----
= -2x^2+24x - 72 + 9x-54 + 1
-----
= -2x^2 + 33x - 125
-------------------
Cheers,
Stan H.